OpenStudy (rayford):

What are the asymptotes of the hyperbola given by the equation x^2/144-y^2/1=1

OpenStudy (amistre64):

slope is the sqrt of y/x denoms

OpenStudy (amistre64):

since the center point is the anchor to the slopes; y = +-mx should suffice

OpenStudy (anonymous):

The asymtopes are y = 1/12x and y = -1/12x. (sorry mods!)

OpenStudy (amistre64):

..... and what would compel you to this life of crime and debauchery? why?!?!?

OpenStudy (rayford):

huh? here are the chpices.y = 6x y =1/36 x y =1/ 6x y = 36x

OpenStudy (amistre64):

none of the them

OpenStudy (amistre64):

wiat .... maybe

OpenStudy (amistre64):

nah, none of them

OpenStudy (amistre64):